a plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for second part of the trip?
Answer by checkley71(8403)

560/(X+1)=480/X
560X=480(X+1)
560X=480X+480
560X-480X=480
80X=480
X=480/80
X=6 HOURS FOR THE SECOND 480 MILE TRIP.
6+1=7 HOURS FOR THE FIRST 560 MILE TRIP.
PROOF
560/7=480/6
80=80